Web Cites Research projects a rate of return of 20 percent on new projects.
Management plans to plow back 30 percent of all earnings into the rm. Earn-
ings this year will be $2 per share, and investors expect a 12 percent rate of
return on the stock.
a) What is the sustainable growth rate?
b) What is the stock price?
c) What is the P/E ratio?
d) What would the price and P/E ratio be if the rm paid out all earnings
as dividends?
e) What do you conclude about the relationship between growth opportuni-
ties and P/E ratios?
f) What is the present value of growth opportunities?
Answer #1

a.

Sustainable Growth Rate = (0.30)(0.12) = 3.60%

b.

Stock Price = 2(0.70)/(0.20 - 0.036)

Stock Price = $8.54

c.

P/E = 8.54/2 = 4.27

d.

Stock Price = 2/(0.20 - 0.036) = $12.20

P/E = 12.20/2 = 6.10